OpenStudy (anonymous):

Please Help! Select the line parallel to 8x-20y=40 a.) y=-4/5x+5 b.) 20x+8y=16 c.) 4x-10y=51 d.) y=5/4x+3

OpenStudy (anonymous):

The answer I get is y=2/5x-20

OpenStudy (anonymous):

I'm unsure how to change the format to fit the options I have...

OpenStudy (asevilla5):

8x-20y=40 so first you want it as y=mx+b so you subtract -8x on both sides to get -20y=-8x+40 then side -20 on both sides. which later becomes y=2/5x-20 it would be c because when you work on that equation you also get 2/5 which is you looking for. they have to have the same mx value

OpenStudy (anonymous):

@Asevilla5 Can you show me how to work the equation to get it to that format?

OpenStudy (asevilla5):

alright for C, its not in y=mx=b so for 4x-10y=51 subtract 4 on both sides to get -10y=-4x+51 then divide by -10 on both sides to get y=2/5+51/-10 for the 51/-10 you can just leave it like that since it is already simplified or you will get a mixed number that you don't want

OpenStudy (anonymous):

Alright that makes more sense =) Thanks I really appreciate it !

OpenStudy (asevilla5):

no problem =)